OpenStudy (anonymous):

Is my answer right? Solve. 1/25 = 5^x + 4 x = 9/2

OpenStudy (anonymous):

@kelly226 please:))))))))

OpenStudy (anonymous):

actually.. we can re-write 1/25 as a power of 2 like this: 1/25 = (25)^(-1) = (5^2)^(-1) = 5^(-1 * 2) = 5^(-2). So, we have: 5^(-2) = 5^(x + 4). and Since the bases are the same, we can compare the exponents to get: -2 = x + 4 ==> x = -6. so, the answer is actually X = -6 :) make sense?

OpenStudy (anonymous):

Wait, how did you get (-1)??

OpenStudy (anonymous):

those were just examples of how u can re write 1/25..

OpenStudy (anonymous):

ohhhh thank you!!!!!!!!!!!!!!!!!!:)))))))))))))))))))))))

OpenStudy (anonymous):

u can just ignore those if u want.. they dont really have anything to do with the problem.. i was just showing u:)

OpenStudy (anonymous):

yw!:)